OpenStudy (anonymous):

Perform the following multiplication. The answer should be in simplified form for radical expressions.

OpenStudy (anonymous):

\[(\sqrt{x}+3)(\sqrt{x}-3)\]

OpenStudy (yttrium):

You remember foil? I saw somebody helping you out about the foil method. Just reply if you still need help :)

OpenStudy (anonymous):

would the square roots cancel out since they're being multiplied together or no?

OpenStudy (yttrium):

Hmm. Let me give you some glimpse about it. Do you know that \[\sqrt{x} = x^\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

No

OpenStudy (anonymous):

So, it just becomes x?

OpenStudy (yttrium):

Yeah. x^(1/2) timex x^(1/2) = x^(1/2 + 1/2) = x

OpenStudy (anonymous):

Well, yea that's what I was asking, but I think I just said it wrong lol

OpenStudy (yttrium):

You are correct. I just give you a proof of why it is correct. :) So, answer it now and let us check if you got it right.

OpenStudy (anonymous):

\[x-3\sqrt{x}+3\sqrt{x}-9?\]

OpenStudy (yttrium):

Yeah. But you can simplify it more.

OpenStudy (anonymous):

the middle terms cancel out? x-9

OpenStudy (yttrium):

Right. Now. Let me give you one more hint about this.

OpenStudy (yttrium):

Whenever you see an algebraic terms like (a+b) (a-b), the answer is always a^2 - b^2

OpenStudy (anonymous):

Oh, i remember that formula!

OpenStudy (yttrium):

now. can you answer this one: simplify (x^2 + 8) (x^2 - 8)

OpenStudy (anonymous):

x^4-64?